# The Open Journal of Astrophysics

### v1

@@ 3.3 Mass slope constraints from standard candle magnifications
Constraining the mass slope $\eta$ of the lens galaxy from photometric observations alone is challenging, particularly for SN Zwicky, which occurred at an offset of 7 kpc from its host galaxy’s centre and therefore lacks strongly lensed host galaxy arcs. This is not an uncommon scenario; around half of strongly lensed SNe is expected to be without a strongly lensed host (Sainz de Murieta et al., 2024). An alternative approach for constraining the mass slope is to utilise the standard candle nature of SNIa, which provides information about the absolute magnification. The magnification of the macro lensing model, $\mu_{\rm th}$, depends strongly on the slope of the mass profile, with flatter profiles providing larger magnifications. However, constraining the mass slope through the observed magnification can only be done when the additional magnification effects of microlensing are taken into consideration. section 4 describes our approach to modelling microlensing effects through simulations, which require inputs for the convergence ($\kappa$) and shear ($\gamma$). We use observations from the Keck $J$-band to model the lens systems of iPTF16geu and SN Zwicky, and determine the $\kappa$ and $\gamma$ values for each SN image, for a range of mass slopes between $\eta=1.4$ and $\eta=2.1$ (step size 0.05). Our range of $\eta$ values was initially chosen to encompass the most common slope values from known lens galaxies, and eventually extended to lower $\eta$ values to capture the $95\%$ confidence level contours for iPTF16geu and SN Zwicky. We also compute the predicted macro magnification $\mu_{\textrm{th}}$ for each SN image and slope value, which, in combination with the simulated microlensing magnification distributions, allows us to determine the best-fit slope values of iPTF16geu and SN Zwicky. Table 1 shows the best-fit convergence, shear and magnification values per slope and per image for SN Zwicky and iPTF16geu, and Fig. 3 shows results and residuals of SN Zwicky’s best-fit macrolens model for a slope of $\eta=1.8$.
@@ 4 Microlensing simulations
The microlensing magnification maps are created using Inverse Polygon Mapping (Mediavilla et al., 2006; Jiménez-Vicente & Mediavilla, 2022) implemented on a GPU for efficiency (Weisenbach et al., in prep). The maps are 20 Einstein radii and 10,000 pixels per side. This is a large enough size to create a representative magnification distribution (Vernardos & Fluke, 2013), while the pixel scale corresponds to approximately 1 day of SN expansion. The maps are convolved with a uniform disk to approximate the size and profile of the expanding SN photosphere at observation, a sufficient approximation for our purposes (Pierel & Rodney, 2019). While more complex SN intensity models may formally be more accurate (Goldstein et al., 2018; Huber et al., 2019), the main source parameter that impacts microlensing is the half-light radius of the source (Mortonson et al., 2005; Vernardos & Tsagkatakis, 2019). For both iPTF16geu and SN Zwicky, the size of the photosphere at the time of observations is of the order of a light day. We average 10 maps for the final microlensing magnification probability distributions in order to minimize scatter from simulating a finite source plane region.
The maps are created using a single mass (in our case, $1M_{\odot}$) for all the microlenses, a common approximation as the magnification distribution is largely insensitive to a mass spectrum, except in extreme scenarios (Schechter et al., 2004). We note that the precise value of the mass used only loosely effects results, as the SN is still only a fraction of an Einstein radius in size, and a smaller microlens mass is degenerate with a larger source size (higher SN expansion speed). The number of microlenses used to create each map depends on the macro-model parameters, but is large enough to ensure that the majority of microlensed flux is accounted for (Katz et al., 1986).
@@ 5.1 Lens galaxy stellar mass estimates
The lens galaxy stellar masses are inferred from a series of photometric observations in the visible and NIR, which are fitted with a spectral energy distribution (SED) model. We use prospector to fit an SED model based on flexible stellar population synthesis (FSPS; Conroy et al., 2009; Conroy & Gunn, 2010b, a; via the python-fsps bindings; Johnson et al., 2024) to our photometric observations of the two lens galaxies. We infer the stellar mass ($M_{*}/M_{\odot}$) jointly with the stellar metallicity ($Z_{*}/Z_{\odot}$), optical depth of dust ($\tau_{5500}$), and two parameters describing the star formation history ($t_{\text{age}}$, $\tau_{\text{SF}}$). Although the photometry alone does not constrain the additional parameters, we leave them free to get a robust estimate of the stellar mass uncertainty. We use the MILES stellar template library (Sánchez-Blázquez et al., 2006), MIST isochrones (Choi et al., 2016), a Chabrier (2003) initial mass function, and a parametric star formation history (SFH) that is linearly increasing at early times and exponentially decaying at late times (see e.g. Carnall et al., 2019; Johnson et al., 2021). The SFH has the form
@@
where $t_{\text{age}}$ is the age of the galaxy, $0<t\leq t_{\text{age}}$ is lookback time, and $t_{\text{age}}-\tau_{\text{SF}}$ sets the peak of the SFH. We use the Calzetti et al. (2000) dust attenuation law, with fixed slope and optical depth $\tau_{5500}$ at 5500 Å. Gas-phase metallicity is coupled to stellar metallicity, with nebular emission being provided by the Cloudy (Ferland et al., 2013) model grids from Byler et al. (2017). We correct for Milky-Way dust using the Gordon et al. (2023) extinction curve (implemented in Gordon, 2024) with $R_{V}=3.1$, and $E(B-V)$ estimated from the Schlafly & Finkbeiner (2011) dust map. We infer $E(B-V)=0.0725$ mag for iPTF16geu, and $E(B-V)=0.1558$ mag for SN Zwicky. We carry out our inference with the lens galaxy redshifts fixed.
The posterior distribution over our five model parameters is sampled using a dynamic nested sampler (Skilling, 2004, 2006; Higson et al., 2019) implemented in dynesty (Speagle, 2020; Koposov et al., 2024), with a proposal strategy based on rejection sampling within ellipsoidal bounds (Feroz et al., 2009). Our priors are: $\log(M_{*}/M_{\odot})\sim\mathcal{U}(8,12)$, $\log(Z_{*}/Z_{\odot})\sim\mathcal{U}(-2.00,0.19)$, $\tau_{5500}\sim\mathcal{U}(0,2)$, $t_{\text{age}}/\text{Gyr}\sim\mathcal{U}(0,t_{\text{Univ}}(z))$, and $\log(\tau_{\text{SF}}/\text{Gyr})\sim\mathcal{U}(-1,\log(30))$. Fig. A.1 in the Appendix shows a corner plot of all the free parameters used in the stellar mass fits for iPTF16geu and SN Zwicky.
Table 2: Photometry and lens fraction measurements for iPTF16geu and SN Zwicky, which serve as input for the stellar mass analysis with prospector. The lens fraction is the ratio between the lens galaxy flux and the total flux. Both the photometry and lens fractions are measured within an aperture with radius $1^{\prime\prime}$. For iPTF16geu, observations in six HST filters are used. For SN Zwicky, HST observations in the $F475W$, $F625W$, and $F814W$ bands, and JWST observations in the $F115W$, $F150W$, $F356W$, and $F444W$ filters are used. We also quote the Milky Way extinction correction (in magnitudes) for each band, computed using the Gordon et al. (2023) curve. The photometric measurements are given in AB magnitudes, and have not been corrected for Milky-Way reddening or the lens fractions.
@@ 5.3 Lens galaxy light model for SN Zwicky
To construct the light model for SN Zwicky’s lens galaxy, we use JWST observations made by NIRCam in the $F115W$ band. The observations were taken one year after the appearance of SN Zwicky, and hence the SN images are not visible anymore. In order to estimate the positions of the SN images on the lens galaxy, we use the HST $F814W$ band observations from when SN Zwicky was active. We combine the observations by projecting the HST data of SN Zwicky on top of the JWST observations using the publicly available software reproject [^4] https://reproject.readthedocs.io/en/stable/ (Robitaille et al., 2020). The projection is done through the World Coordinate Systems of both observations. We correct remaining shifts between the observations by fitting five stars in the field with a Moffat profile and correcting for their offsets. To ensure the clearest representation of SN Zwicky’s image positions, we use the difference image between HST’s observations with the active SN and the reference observations taken a year later. Fig. 1 shows the JWST data with the projected SN Zwicky image positions. Finally, we fit a Moffat profile to the projected SN images to accurately determine their positions within the JWST observation.
@@ 6.1 Observed quantities
  • Observed total magnification $\bm{\hat{\mu}}$: the observed magnification is determined by comparing the observed apparent magnitudes of the lensed SNe with the expected magnitude of a standard SNIa at that redshift. We take the results for the total magnification (sum of all images) from Dhawan et al. (2020) for iPTF16geu, $\hat{\mu}=67.8\pm 2.9$, and from Goobar et al. (2023) for SN Zwicky, $\hat{\mu}=23.7\pm 3.2$.
@@
  • SN image flux ratios $\bm{\hat{R}_{i}}$: the observed flux ratios of the SN images, normalised to one. We take the results from Dhawan et al. (2020) for iPTF16geu, $\hat{R_{i}}=[0.52,0.23,0.11,0.13]$, and from Goobar et al. (2023) for SN Zwicky, $\hat{R_{i}}=[0.37,0.15,0.30,0.18]$.
@@ 6.3 Derived parameters
    with $G$ the gravitational constant, $c$ the speed of light, and $D_{\rm l}$, $D_{\rm s}$ and $D_{\rm ls}$ the angular diameter distances between the observer and the lens, the observer and the source, and the lens and the source, respectively. Distances are calculated using a standard flat $\Lambda$CDM model with $H_{0}=67.4\ \textrm{km}\,\textrm{s}^{-1}\textrm{Mpc}^{-1}$ and $\Omega_{\textrm{m}}=0.315$ (Planck Collaboration et al., 2020). Finally, the stellar mass fraction is given by
@@ 6.5 Likelihoods on data
  • $\bm{P(\hat{\mu}\,|\,\mu)}$. The observed total magnification, $\hat{\mu}$, is determined by comparing the apparent magnitudes of the lensed SNe with the expected magnitudes of standard SNe Ia at the same redshifts. For iPTF16geu, we adopt the value $\hat{\mu}=67.8\pm 2.9$ from Dhawan et al. (2020). For SN Zwicky, we use $\hat{\mu}=23.7\pm 3.2$ from Goobar et al. (2023). The likelihood $P(\hat{\mu}\,|\,\mu)$ is modelled as a Gaussian distribution with mean and standard deviation equal to the reported values in the respective studies.
@@ 6.6 Priors
[…]
@@
[…]
@@ 7.1 Stellar mass fractions
The stellar mass fraction of the lens galaxy at the SN image positions, $f_{*,i}$, is a derived parameter in our model, calculated from the stellar mass, $M_{*}$, the lens galaxy flux, $\hat{F}_{\textrm{lens}}$, the lens galaxy flux density at the SN image positions, $\hat{S}_{\textrm{lens},i}$, and the convergence at the image positions, $\kappa_{i}$. Fig. 8 shows the inferred stellar mass fractions per lensed SN image, for iPTF16geu and SN Zwicky. Both iPTF16geu and SN Zwicky have high stellar mass fractions, which can be explained by the fact that the SN images are formed close to the lens centre, where the stellar density is high compared to the dark matter density. The median $f_{*,i}$ values for iPTF16geu ($\sim 0.7-0.8$) are higher than the ones inferred for iPTF16geu in Mörtsell et al. (2020) ($\sim 0.2$), which is due to differences in the stellar mass determination between the papers. In Mörtsell et al. (2020), the stellar mass was inferred from scaling relations in Kauffmann et al. (2003), whereas in this work, we model the full lens galaxy SED with prospector.
@@ 7.2 Dark compact objects and galaxy mass slopes
Fig. 9 shows the joint $\eta$ and $f_{\rm dc}$ posterior distributions for iPTF16geu and SN Zwicky. For both lensed SNe, there is no preference for additional dark compact objects, and stars alone are sufficient to explain the observed microlensing effects. We put upper limits at the $95\%$ confidence level on the fraction of dark compact objects of $f_{\rm dc}$ $<0.25$ for iPTF16geu and $f_{\rm dc}$ $<0.47$ for SN Zwicky. The upper limits from iPTF16geu are tighter than those from SN Zwicky, due to the fact that iPTF16geu has high stellar mass fractions ($\sim 0.7-0.8$), leaving less room for the presence of dark compact objects. These limits apply to compact objects with masses greater than $0.02M_{\odot}$. Microlenses with lower masses have Einstein radii smaller than the size of the SN photosphere, causing their magnification contributions to average out.
By utilising the standard candle nature of SNIa, while incorporating microlensing into our modelling, we were able to constrain the galaxy mass slopes for these systems, which are difficult to infer from lens modelling alone. For iPTF16geu, we find a slope of $\eta=1.70\pm 0.07$, which is slightly lower but consistent with the value derived in Mörtsell et al. (2020) ($\eta\sim 1.8$). SN Zwicky’s mass slope is $\eta=1.81\pm 0.10$. The step-like features in the posterior distribution of $\eta$ reflect the discreteness introduced by the microlensing simulations.
@@
[…]
For both iPTF16geu and SN Zwicky, the brightest observed images are $\sim 0.7$ magnitudes brighter due to microlensing, which is in line with microlensing expectations for lensed SNe (Dobler & Keeton, 2006), especially considering those images are both saddle points, which are more strongly affected by microlensing.
In Fig. 10, we plot our inferred mass slopes for iPTF16geu and SN Zwicky together with those from lens galaxies in the Sloan Lens ACS Survey (SLACS; Auger et al., 2010) and in the Strong Lensing Legacy Survey (SL2S; Sonnenfeld et al., 2013), as a function of redshift and stellar mass. iPTF16geu and SN Zwicky’s mass slopes are relatively low compared to the lens galaxy population, while still being within the range observed for other galaxies. In terms of stellar mass, SN Zwicky is significantly lighter than the other lens galaxies. As was pointed out in Goobar et al. (2023), iPTF16geu and SN Zwicky are the first discoveries of a new class of extremely low-mass and compact lens systems, identified primarily due to their high magnification. These discoveries were only possible because the lensed sources were SNIa with standardisable brightnesses. In this work, we have demonstrated that these lensing systems are highly suitable for microlensing studies and for constraining the matter in compact objects.

### v2

@@ 3.4 Mass slope constraints from standard candle magnifications
Constraining the mass slope $\eta$ of the lens galaxy from photometric observations alone is challenging, particularly for SN Zwicky, which occurred at an offset of 7 kpc from its host galaxy’s centre and therefore lacks strongly lensed host galaxy arcs. This is not an uncommon scenario; around half of strongly lensed SNe is expected to be without a strongly lensed host (SainzdeMurieta2024). An alternative approach for constraining the mass slope is to utilise the standard candle nature of SNIa, which provides information about the absolute magnification. The magnification of the macro lensing model, $\mu_{\rm th}$, depends strongly on the slope of the mass profile, with flatter profiles providing larger magnifications. However, constraining the mass slope through the observed magnification can only be done when the additional magnification effects of microlensing are taken into consideration. Section 4 describes our approach to modelling microlensing effects through simulations, which require inputs for the convergence ($\kappa$) and shear ($\gamma$). We use observations from the Keck $J$-band to model the lens systems of iPTF16geu and SN Zwicky, and determine the $\kappa$ and $\gamma$ values for each SN image, for a range of mass slopes between $\eta=1.4$ and $\eta=2.1$ (step size 0.05). Our range of $\eta$ values was initially chosen to encompass the most common slope values from known lens galaxies, and eventually extended to lower $\eta$ values to capture the $95\%$ confidence level contours for iPTF16geu and SN Zwicky. We also compute the predicted macro magnification $\mu_{\textrm{th}}$ for each SN image and slope value, which, in combination with the simulated microlensing magnification distributions, allows us to determine the best-fit slope values of iPTF16geu and SN Zwicky. Table 1 shows the best-fit convergence, shear and magnification values per slope and per image for SN Zwicky and iPTF16geu, and Fig. 3 shows results and residuals of SN Zwicky’s best-fit macrolens model for a slope of $\eta=1.8$.
@@ 4 Microlensing simulations
The microlensing magnification maps are created using Inverse Polygon Mapping (2006ApJ...653..942M; 2022ApJ...941...80J) implemented on a GPU for efficiency (2025MNRAS.541..281W). The maps are 20 Einstein radii and 10,000 pixels per side. This is a large enough size to create a representative magnification distribution (2013MNRAS.434..832V), while the pixel scale corresponds to approximately 1 day of SN expansion. The maps are convolved with a uniform disk to approximate the size and profile of the expanding SN photosphere at observation, a sufficient approximation for our purposes (pierel2019turning). While more complex SN intensity models may formally be more accurate (Goldstein2018microlensing; Huber2019LSSTcadence), the main source parameter that impacts microlensing is the half-light radius of the source (2005ApJ...628..594M; Vernardos2019). For both iPTF16geu and SN Zwicky, the size of the photosphere at the time of observations is of the order of a light day. We average 10 maps for the final microlensing magnification probability distributions in order to minimize scatter from simulating a finite source plane region.
The maps are created using a single mass (in our case, $1M_{\odot}$) for all the microlenses, a common approximation as the magnification distribution is largely insensitive to a mass spectrum, except in extreme scenarios (2004ApJ...613...77S). We note that the precise value of the mass used only loosely effects results, as the SN is still only a fraction of an Einstein radius in size, and a smaller microlens mass is degenerate with a larger source size (higher SN expansion speed). The number of microlenses used to create each map depends on the macro-model parameters, but is large enough to ensure that the majority of microlensed flux is accounted for (1986ApJ...306....2K).
@@ 5.1 Lens galaxy stellar mass estimates
The lens galaxy stellar masses are inferred from a series of photometric observations in the visible and NIR, which are fitted with a spectral energy distribution (SED) model. We use prospector to fit an SED model based on flexible stellar population synthesis (FSPS; Conroy2009; Conroy2010; Conroy2010ApJ; via the python-fsps bindings; Johnson2024) to our photometric observations of the two lens galaxies. We infer the stellar mass ($M_{*}/M_{\odot}$) jointly with the stellar metallicity ($Z_{*}/Z_{\odot}$), optical depth of dust ($\tau_{5500}$), and two parameters describing the star formation history ($t_{\text{age}}$, $\tau_{\text{SF}}$). Although the photometry alone does not constrain the additional parameters, we leave them free to get a robust estimate of the stellar mass uncertainty. We use the MILES stellar template library (SanchezBlazquez2006), MIST isochrones (Choi2016), a Chabrier2003 initial mass function, and a parametric star formation history (SFH) that is linearly increasing at early times and exponentially decaying at late times (see e.g. Carnall2019; Johnson2021). The SFH has the form
@@
where $t_{\text{age}}$ is the age of the galaxy, $0<t\leq t_{\text{age}}$ is lookback time, and $t_{\text{age}}-\tau_{\text{SF}}$ sets the peak of the SFH. We use the Calzetti2000 dust attenuation law, with fixed slope and optical depth $\tau_{5500}$ at 5500 Å. Gas-phase metallicity is coupled to stellar metallicity, with nebular emission being provided by the Cloudy (Ferland2013) model grids from Byler2017. We correct for Milky-Way dust using the Gordon2023 extinction curve (implemented in Gordon2024, and relying on work from Gordon2009; Fitzpatrick2019; Gordon2021; Decleir2022) with $R_{V}=3.1$, and E(BV)E(B-V) estimated from the Schlafly2011 dust map. We infer E(BV)=0.0725E(B-V)=0.0725 mag for iPTF16geu, and E(BV)=0.1558E(B-V)=0.1558 mag for SN Zwicky. We carry out our inference with the lens galaxy redshifts fixed.
The posterior distribution over our five model parameters is sampled using a dynamic nested sampler (Skilling2004; Skilling2006; Higson2019) implemented in dynesty (Speagle2020; Koposov2024), with a proposal strategy based on rejection sampling within ellipsoidal bounds (Feroz2009). Our priors are: $\log(M_{*}/M_{\odot})\sim\mathcal{U}(8,12)$, $\log(Z_{*}/Z_{\odot})\sim\mathcal{U}(-2.00,0.19)$, $\tau_{5500}\sim\mathcal{U}(0,2)$, $t_{\text{age}}/\text{Gyr}\sim\mathcal{U}(0,t_{\text{Univ}}(z))$, and $\log(\tau_{\text{SF}}/\text{Gyr})\sim\mathcal{U}(-1,\log(30))$. Fig. A.1 in the Appendix shows a corner plot of all the free parameters used in the stellar mass fits for iPTF16geu and SN Zwicky.
Table 2: Photometry and lens fraction measurements for iPTF16geu and SN Zwicky, which serve as input for the stellar mass analysis with prospector. The lens fraction is the ratio between the lens galaxy flux and the total flux. Both the photometry and lens fractions are measured within an aperture with radius $1^{\prime\prime}$. For iPTF16geu, observations in six HST filters are used. For SN Zwicky, HST observations in the $F475W$, $F625W$, and $F814W$ bands, and JWST observations in the $F115W$, $F150W$, $F356W$, and $F444W$ filters are used. We also quote the Milky Way extinction correction (in magnitudes) for each band, computed using the Gordon2023 curve. The photometric measurements are given in AB magnitudes, and have not been corrected for Milky-Way reddening or the lens fractions.
@@ 5.3 Lens galaxy light model for SN Zwicky
To construct the light model for SN Zwicky’s lens galaxy, we use JWST observations made by NIRCam in the $F115W$ band. The observations were taken one year after the appearance of SN Zwicky, and hence the SN images are not visible anymore. In order to estimate the positions of the SN images on the lens galaxy, we use the HST $F814W$ band observations from when SN Zwicky was active. We combine the observations by projecting the HST data of SN Zwicky on top of the JWST observations using the publicly available software reproject [^4] https://reproject.readthedocs.io/en/stable/ (Reproject2020). The projection is done through the World Coordinate Systems of both observations. We correct remaining shifts between the observations by fitting five stars in the field with a Moffat profile and correcting for their offsets. To ensure the clearest representation of SN Zwicky’s image positions, we use the difference image between HST’s observations with the active SN and the reference observations taken a year later. Fig. 1 shows the JWST data with the projected SN Zwicky image positions. Finally, we fit a Moffat profile to the projected SN images to accurately determine their positions within the JWST observation.
@@ 6.1 Observed quantities
  • Observed total magnification $\bm{\hat{\mu}}$: the observed magnification is determined by comparing the observed apparent magnitudes of the lensed SNe with the expected magnitude of a standard SNIa at that redshift. We take the results for the total magnification (sum of all images) from Dhawan2020_16geu for iPTF16geu, $\hat{\mu}=67.8\pm 2.9$, and from Goobar2023_SNZwicky for SN Zwicky, $\hat{\mu}=23.7\pm 3.2$.
@@
  • SN image flux ratios $\bm{\hat{R}_{i}}$: the observed flux ratios of the SN images, normalised to one. We take the results from Dhawan2020_16geu for iPTF16geu, $\hat{R_{i}}=[0.52,0.23,0.11,0.13]$, and from Goobar2023_SNZwicky for SN Zwicky, $\hat{R_{i}}=[0.37,0.15,0.30,0.18]$.
@@ 6.3 Derived parameters
[…]
@@ 6.5 Likelihoods on data
  • $\bm{P(\hat{\mu}\,|\,\mu)}$. The observed total magnification, $\hat{\mu}$, is determined by comparing the apparent magnitudes of the lensed SNe with the expected magnitudes of standard SNe Ia at the same redshifts. For iPTF16geu, we adopt the value $\hat{\mu}=67.8\pm 2.9$ from Dhawan2020_16geu. For SN Zwicky, we use $\hat{\mu}=23.7\pm 3.2$ from Goobar2023_SNZwicky. The likelihood $P(\hat{\mu}\,|\,\mu)$ is modelled as a Gaussian distribution with mean and standard deviation equal to the reported values in the respective studies.
@@ 6.6 Priors
where $P(\mu\hat{R}_{i}\,|\,M_{*},\eta,f_{\rm dc},\hat{F}_{\textrm{lens}},\hat{S}_{\textrm{lens},i})$ is the microlensing magnification distribution for image $i$ evaluated at $\mu_{i}=\mu\hat{R}_{i}$. In practice, the total magnification $\mu$ is sampled from a uniform proposal distribution, $\mathcal{U}(\mu\,;\,15,50)$, and multiplied by the SN image flux ratios $\hat{R}_{i}$ to obtain $\mu_{i}$. For the remaining free parameters, we use the following priors:
@@
$\prod_{i=1}^{N}\mathcal{U}(f_{\rm dc}+f_{*,i}\,;\,0,1).$ (12)
The prior in eq. 12 ensures that for each SN image, the fraction of compact objects $f_{c,i}=f_{\rm dc}+f_{*,i}$ is between 0 and $1$. For $P(M_{*})$ and $P(\eta)$, we initialise the walkers in a range of $\mathcal{U}(M_{*}\,;\,10^{9},10^{10.8})$ and $\mathcal{U}(\eta\,;\,1.4,2.1)$, and we have uninformative priors over the entire parameter space. The values of $\eta$ are sampled from a continuous distribution but are rounded to the nearest $0.05$ to match the precision used in the microlensing simulations. Similarly, the compact object fraction per image, $f_{\textrm{c},i}=f_{\textrm{dc}}+f_{*,i}$ is rounded to the nearest $0.1$ to align with the simulation’s precision.
@@ 7.1 Stellar mass fractions
The stellar mass fraction of the lens galaxy at the SN image positions, $f_{*,i}$, is a derived parameter in our model, calculated from the stellar mass, $M_{*}$, the lens galaxy flux, $\hat{F}_{\textrm{lens}}$, the lens galaxy flux density at the SN image positions, $\hat{S}_{\textrm{lens},i}$, and the convergence at the image positions, $\kappa_{i}$. Fig. 8 shows the inferred stellar mass fractions per lensed SN image, for iPTF16geu and SN Zwicky. Both iPTF16geu and SN Zwicky have high stellar mass fractions, which can be explained by the fact that the SN images are formed close to the lens centre, where the stellar density is high compared to the dark matter density. The median $f_{*,i}$ values for iPTF16geu ($\sim 0.7-0.8$) are higher than the ones inferred for iPTF16geu in Mortsell2020 ($\sim 0.2$), which is due to differences in the stellar mass determination between the papers. In Mortsell2020, the stellar mass was inferred from scaling relations in Kauffmann2003, whereas in this work, we model the full lens galaxy SED with prospector.
@@ 7.2 Dark compact objects and galaxy mass slopes
Fig. 9 shows the joint $\eta$ and $f_{\rm dc}$ posterior distributions for iPTF16geu and SN Zwicky. For both lensed SNe, there is no preference for additional dark compact objects, and stars alone are sufficient to explain the observed microlensing effects. We put upper limits at the $95\%$ confidence level on the fraction of dark compact objects of $f_{\rm dc}$ <0.25<0.25 for iPTF16geu and $f_{\rm dc}$ <0.47<0.47 for SN Zwicky. The upper limits from iPTF16geu are tighter than those from SN Zwicky, due to the fact that iPTF16geu has high stellar mass fractions ($\sim 0.7-0.8$), leaving less room for the presence of dark compact objects. These limits apply to compact objects with masses greater than $0.02M_{\odot}$. Microlenses with lower masses have Einstein radii smaller than the size of the SN photosphere, causing their magnification contributions to average out.
By utilising the standard candle nature of SNIa, while incorporating microlensing into our modelling, we were able to constrain the galaxy mass slopes for these systems, which are difficult to infer from lens modelling alone. For iPTF16geu, we find a slope of $\eta=1.70\pm 0.07$, which is slightly lower but consistent with the value derived in Mortsell2020 ($\eta\sim 1.8$). SN Zwicky’s mass slope is $\eta=1.81\pm 0.10$. The step-like features in the posterior distribution of $\eta$ reflect the discreteness introduced by the microlensing simulations.
@@
$\textrm{iPTF16geu: }-0.72,\ -0.05,\ 0.78,\ 0.97\ \textrm{mag}.$
$\textrm{SN Zwicky: }-0.68,\ 0.79,\ 0.10,\ 0.74\ \textrm{mag}.$
For both iPTF16geu and SN Zwicky, the brightest observed images are $\sim 0.7$ magnitudes brighter due to microlensing, which is in line with microlensing expectations for lensed SNe (dobler2006microlensing), especially considering those images are both saddle points, which are more strongly affected by microlensing.
In Fig. 10, we plot our inferred mass slopes for iPTF16geu and SN Zwicky together with those from lens galaxies in the Sloan Lens ACS Survey (SLACS; Auger2010) and in the Strong Lensing Legacy Survey (SL2S; Sonnenfeld2013), as a function of redshift and stellar mass. iPTF16geu and SN Zwicky’s mass slopes are relatively low compared to the lens galaxy population, while still being within the range observed for other galaxies. In terms of stellar mass, SN Zwicky is significantly lighter than the other lens galaxies. As was pointed out in Goobar2023_SNZwicky, iPTF16geu and SN Zwicky are the first discoveries of a new class of extremely low-mass and compact lens systems, identified primarily due to their high magnification. These discoveries were only possible because the lensed sources were SNIa with standardisable brightnesses. In this work, we have demonstrated that these lensing systems are highly suitable for microlensing studies and for constraining the matter in compact objects.